OpenStudy (anonymous):

Find both the x-intercept and the y-intercept of the line given by the equation ; 9.4x-5.8y-8.3=0 . Round your answers to 2 decimal points. x-intercept = ? and y-intercept = ?

OpenStudy (anonymous):

At the X-intercept, y = 0 At the Y- intercept, x = 0 To find x - intercept, substitue y = 0 into the equation 9.4x-5.8y-8.3=0 9.4x-5.8(0)-8.3=0 9.4x-0-8.3=0 9.4x=8.3 x=8.3/9.4 X=0.88 Likewise, to find y-intercept, subst. X = 0 into the equation too 9.4(0)-5.8y-8.3=0 0-5.8y-8.3=0 5.8y=-8.3 Y=-8.3/5.8 Y=-1.43 x-intercept: 0.88 y-intercept: 1.43 Anyway, hope that i helped. :)

OpenStudy (anonymous):

I got .88! but somehow I got y=2.5 :/

OpenStudy (anonymous):

if it says round to 2 decimal points, is .88 correct?

OpenStudy (anonymous):

I got 1.43 now, thank you.

OpenStudy (anonymous):

Sorry. It's -1.43 and you are welcome. ^_^

OpenStudy (anonymous):

ok i didn't get -1.43 ... lol

OpenStudy (anonymous):

Study the solution i provided again. :/ That's the only solution that i can provide. I have no idea why you got a positive answer unless you show me your steps.